OpenStudy (anonymous):

Carl takes a random sample of 100 students in his school and finds that 72% of the sample prefers pizza over burgers. There are 1,500 students in the school. Based on the sample proportion, how many students in the school would be expected to prefer pizza over burgers?

OpenStudy (anonymous):

@Michele_Laino help ??????????????????????????

OpenStudy (anonymous):

Carl takes a random sample of 100 students in his school and finds that 72% of the sample prefers pizza over burgers. There are 1,500 students in the school. Based on the sample proportion, how many students in the school would be expected to prefer pizza over burgers? 72 100 1,080 1,500

OpenStudy (michele_laino):

if 72 students (every 100 students) prefer pizza over burgers, then the number of students (every 1500 students in total) which prefer pizza over burgers, will be: \[\Large \frac{{1500}}{{100}} \times 72 = ...?\]

OpenStudy (anonymous):

1,080

OpenStudy (michele_laino):

if the sample of students is enlarged, the proportion has to remain the same

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

ok i have 3 more questions is that ok ?

OpenStudy (anonymous):

Ben takes a random sample of 25 students in his seventh-grade class and finds that 85% of the sample prefers math over science. There are 150 students in the seventh grade. Based on the sample proportion, about how many students in the seventh grade would be expected to prefer math over science? 25 85 150 128

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

ok!

OpenStudy (anonymous):

Ben takes a random sample of 25 students in his seventh-grade class and finds that 85% of the sample prefers math over science. There are 150 students in the seventh grade. Based on the sample proportion, about how many students in the seventh grade would be expected to prefer math over science? 25 85 150 128

OpenStudy (michele_laino):

here, every 100 students, 85 prefer maths over science, so, as in the previous exercise, we have to compute this quantity: \[\Large \frac{{150}}{{100}} \times 85 = ...?\]

OpenStudy (anonymous):

127.5

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

you have to round off, that result, so what you get?

OpenStudy (anonymous):

128

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

Angela wants to know how many families in her neighborhood plan to attend the parade. She puts all 120 of the neighborhood addresses in a hat and draws a random sample of 30 addresses. She then asks those families if they plan to attend the parade. She finds that 40% of the families plan to attend the parade. She claims that 40% of the neighborhood families would be expected to attend the parade. Is this a valid inference? Yes, this is a valid inference because the 30 families speak for the whole neighborhood Yes, this is a valid inference because she took a random sample of the neighborhood No, this is not a valid inference because she did not take a random sample of the neighborhood No, this is not a valid inference because she asked only 30 families

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

please wait a moment

OpenStudy (michele_laino):

hint: a function f(t) is periodic with period T, if checks this condition: \[f\left( {t + T} \right) = f\left( t \right)\]

OpenStudy (michele_laino):

sorry I have wrote the answer of another question @Marc1313

OpenStudy (anonymous):

ok what?

OpenStudy (michele_laino):

I think it is the last option, since Angela should ask to all of 120 families

OpenStudy (anonymous):

ok that you so much!

OpenStudy (michele_laino):

the sample, which Angela has to consider has to be as large as possible

OpenStudy (michele_laino):

thank you!